OpenStudy (anonymous):

the line through (6,-2) with slope -3/4 in point-slope form

OpenStudy (anonymous):

y-y1 = m(x-x1) y +2 = -0.75(x-6) y = -0.75x+6*0.75 -2 simply after that

OpenStudy (anonymous):

*simplify

OpenStudy (anonymous):

@Sheng , if you simplify that, then it will no longer be in point-slope form.

OpenStudy (anonymous):

eqn is4y+3x=10 divide by the coefficient of yand write it as y=-3/4x+10/4